Question 9: An ideal gas has average temperature of 40 °C. What is the average kinetic energy of the gas?
Answer #1

Temperature, T=40oC= 313K

Therefore, the average kinetic energy of the gas per molecule is,

E=\frac{3}{2}\frac{R}{N}T

Where R is the gas constant and N is the Avogrdro's number.

Therefore, 3 8.314 E=of 6.023 + 1023 四*313 = 6.48 + 10-21]